OpenStudy (anonymous):

Which of the following is the solution of :

OpenStudy (anonymous):

\[\log_{x+3} 0.001= -3\]

OpenStudy (anonymous):

A. x=5, B. x=7, C. x=13, D. x=15

OpenStudy (gorv):

@nomad96 is it x+3 ??

OpenStudy (anonymous):

yes

OpenStudy (gorv):

x=7

OpenStudy (gorv):

got it or not

OpenStudy (anonymous):

(x+3)^(-3)=0,001 (x+3)^(-3)=10^(-3) x+3=10 x=7 and @gorv is right :)

OpenStudy (anonymous):

Awe thanks guys!!

OpenStudy (anonymous):

welcome :)